↓ ↓ ↓ ↓ SUBSCRIBE:↓ ↓ ↓ Follow by Email↓ ↓ ↓ ↓

Wednesday, April 25, 2018

Some adjustments

Derek made some last adjustments to the Acme projector today. At first, I figured that it worked better but still didn't work. However, I then tried putting it into the "rewind" mode which always boosted the power for some reason. Low and behold it started to run great. We watched a bit of the trailer for Robocop. Later I dropped it at Any Electronics for a new power plug to be installed as the existing one needed to be jerry-rigged to work and it doesn't look or feel safe (although it mostly is). 
William got a hair cut. Looks pretty good. Thanks Gerald. 







No comments: